# Out-of-Distribution Semantic Occupancy Prediction

### v1

@@ 4.2 Overall Architecture
The OccOoD framework, illustrated in Fig. 3, integrates front-view and BEV representations for precise 3D semantic occupancy prediction. While 3D voxels preserve fine-grained geometric details, BEV enhances spatial reasoning and global scene understanding. To effectively merge these representations, we propose Voxel-BEV Progressive Fusion (VBPF), which iteratively refines voxel-BEV integration, leveraging voxel geometry and BEV efficiency for improved feature representation. This multi-perspective fusion enhances generalization and robustness, significantly boosting out-of-distribution detection performance. Initially, following SGN [66], an image encoder extracts 2D features from RGB images, which are then mapped into 3D space using a view transformation module adapted from MonoScene [12]. The coarse 3D features are refined through geometry guidance, and an auxiliary 3D occupancy head is optimized using binary cross-entropy loss:
$\small L_{\text{geo}}=-\sum_{i}\left[(1-Y_{o,i})\log(1-\hat{Y}_{o,i})+Y_{o,i}\log(\hat{Y}_{o,i})\right],$ (7)
where $i$ indexes voxels, $Y_{o}$ represents ground truth occupancy, and $\hat{Y}_{o}$ indicates the predicted occupancy.
After the proposal network filters the 3D features, optimized with a binary cross-entropy loss $L_{\text{occ}}$, the filtered features undergo Geometry Parsing and Semantic Refinement to extract multi-scale BEV features and fine-grained semantic voxel features. These BEV features are fused with the original 3D features through a UNet-like module, ensuring precise alignment and enhanced representation learning. Finally, the BEV representation is upsampled, concatenated with the aggregated voxel representation, and refined using convolutions to produce the semantic occupancy prediction.
4.3 Voxel-BEV Progressive Fusion
The coarse voxel features $F^{3D}$ from the view transformation module and the seed features $F_{\text{s,0}}^{3D}$ obtained through the proposal network are both 3D voxel features. However, a single representation limits fine-grained semantics and segmentation boundaries, hindering OoD detection. To overcome this, VBPF progressively integrates voxel-based and BEV representations, leveraging BEV’s top-down perspective for global context and spatial uncertainty identification.
Geometry Parsing: The coarse voxel features $F^{3D}$ are processed through a geometric parsing branch, comprising a linear layer, residual connections, a BEV projection module, and a Vision-RWKV (VRWKV) module [26] (see Fig. 3), to encode rich geometric information into compact BEV representations.
Specifically, $F^{3D}$ is initially projected into a lower-dimensional space via a linear layer, refined through a residual block, aligned along the z-axis, and compressed using 2D convolutions to generate dense BEV features. These BEV features are further refined by the VRWKV module, which enhances their representational quality through spatial and channel mixing operations, producing dense geometric BEV features $\{F_{\text{g}_{i}}^{\text{BEV}}\}_{i}$.
@@ 4.3 Voxel-BEV Progressive Fusion
Semantic Refinement: The seed voxels $F_{\text{s,0}}^{3D}$ generated by the proposal network undergo semantic refinement to enhance semantic accuracy. The semantic refinement module comprises three cascaded sparse encoder blocks (SEB), each linked with a BEV projection module and a VRWKV block. Each SEB contains a Sparse Feature Encoder (SFE) [107] and a Sparse Geometry Feature Encoder (SGFE) [107], which optimize the initial semantic representation and refine it through hierarchical pooling and attention mechanisms, capturing fine-grained local details and broader contextual dependencies.
First, the voxel features pass through the SEB, where two SEBs output contextual features at different scales, $F_{\text{s,1}}^{3D}$ and $F_{\text{s,2}}^{3D}$, each embedding rich geometric and contextual information. These features are projected into BEV space via a BEV projection operation, where max-pooling aggregates features with the same BEV index, generating sparse BEV features, which are then converted into dense semantic-BEV features using Spconv’s densification function.
The VRWKV module is then applied to model long-range dependencies, enhancing the representational quality of the BEV features. The final output includes multi-scale semantic BEV features: $\{F_{\text{s}_{i}}^{\text{BEV}}\}_{i}$, as well as the refined semantic feature $F_{\text{s}}^{3D}$ obtained by Eq. (8):
@@
The semantic refinement stage is optimized using a combination of cross-entropy loss and Lovasz loss [6]:
@@
where $Y_{s}$ represents the semantic labels of seed voxels indexed from the semantic scene labels $Y$.
Cross-View Feature Synergy: To effectively integrate BEV and voxel-level representations, we first extract features independently. For BEV processing, we employ a UNet-like architecture, where the encoder captures hierarchical spatial features and the decoder reconstructs high-resolution feature maps for seamless fusion. The inputs include the projection of 3D features into BEV space ($F^{\mathrm{BEV}}_{c}$), geometric BEV features $\{F_{\text{g}_{i}}^{\text{BEV}}\}_{i}$ from the first stage, and semantic BEV features $\{F_{\text{s}_{i}}^{\text{BEV}}\}_{i}$ from the second stage.
The BEV encoder downsamples $F^{\mathrm{BEV}}_{c}$ through convolutional blocks, each followed by a VRWKV block, and integrates geometric BEV features $F_{\text{g}_{i}}^{\text{BEV}}$ and semantic BEV features $F_{\text{s}_{i}}^{\text{BEV}}$ using an ARF module [65]. The process can be represented as Eq. (10):
@@
where $F^{\text{BEV}}_{k_{i}}$ and $A^{\text{BEV}}_{k_{i}}$ represent the $k$-th class feature map of the $i$-th BEV feature and its corresponding attention weight, respectively. The decoder restores spatial resolution using transposed convolutions and skip connections. The processed BEV features are reshaped and upsampled to generate a voxel-level representation $\hat{Y}_{\text{$b$}}$.
For voxel-level feature processing, we adopt the approach of SGN [66] for voxel handling. Specifically, we fuse and refine three types of features: $F_{\text{s}}^{3D}$ obtained from semantic refinement, non-seed voxel features $F_{\text{n}}^{3D}$ filtered by the proposal network, and $F^{3D}_{o}$ from the proposal network. This procedure is formally expressed in Eq. (11):
$\hat{Y}_{\text{$v$}}=\text{MSSP}([F_{\text{s}}^{3D},F_{\text{n}}^{3D},F_{\text{$o$}}^{3D}]),$ (11)
where MSSP denotes the multi-scale semantic propagation.
Subsequently, $\hat{Y}_{\text{$v$}}$ is concatenated with $\hat{Y}_{\text{$b$}}$ and processed through a 1D convolutional layer to generate the final unified representation $\hat{Y}_{f}$. We optimize the model using a combination of scene- and class-wise semantic loss $L_{sem}^{scal}$, geometry $L_{geo}^{scal}$, and cross-entropy loss $L_{ce}$. The overall loss function is defined in Eq. (12):
@@
Loss Function: The total training loss is defined as follows:
@@ 4.4 OoD Occupancy Prediction
Geometry Prior: To more efficiently focus on non-empty regions, we introduce a geometry prior by applying an argmax operation to $\hat{Y}_{f}$. This strategy prioritizes potentially anomalous non-empty regions, significantly improving the effectiveness and accuracy of OoD detection.
VAA-KITTI VAA-KITTI-360
[…]
0.8m 1.0m 1.2m 0.8m 1.0m 1.2m
SGN-S [66] 16.00 24.08 35.09 65.99 17.42 19.99 23.51 66.31
OccOoD-S (ours) 19.64 27.75 38.69 65.91 22.89 25.59 29.21 67.34
Table 1: Quantitative results across different region ranges on the VAA-KITTI and VAA-KITTI-360 datasets. The best results are in bold.
@@
VoxFormer [51] 43.21 54.10 26.90 25.10 7.30 23.50 21.70 3.60 1.90 1.60 4.10 24.40 8.10 24.20 1.60 1.10 0.00 13.10 6.60 5.70 13.41
SGN [66] 45.42 60.40 31.40 28.90 8.70 28.40 25.40 4.50 0.90 1.60 3.70 27.40 12.60 28.40 0.50 0.30 0.10 18.10 10.00 8.30 15.76
Symphonies [41] 42.19 58.40 29.30 26.90 11.70 24.70 23.60 3.20 3.60 2.60 5.60 24.20 10.00 23.10 3.20 1.90 2.00 16.10 7.70 8.00 15.04
CGFormer [110] 44.41 64.30 34.20 34.10 12.10 25.80 26.10 4.30 3.70 1.30 2.70 24.50 11.20 29.30 1.70 3.60 0.40 18.70 8.70 9.30 16.63
OccOoD (Ours) 45.11 61.80 31.60 30.50 10.00 27.90 26.60 2.60 1.00 0.80 5.60 26.70 11.90 28.90 0.50 0.80 0.00 18.70 9.40 5.70 15.85
@@
SGN-S [66] 39.07 46.62 26.54 15.88 0.12 18.34 20.51 0.62 0.44 0.35 1.23 24.76 6.40 21.78 0.48 1.29 0.00 12.25 8.70 4.57 11.10
OccOoD-S (Ours) 40.04 50.17 27.42 18.4 0.01 17.38 20.19 0.74 0.26 0.56 1.04 25.13 5.53 23.56 0.67 1.95 0.00 11.35 8.15 5.80 11.49
Table 2: Quantitative results on the SemanticKITTI test set [3] and VAA-KITTI. The best and the second best results are in bold and underlined, respectively.
Semantic-aware Anomaly Scoring: We categorize voxels into two groups based on their semantic predictions: instance (e.g., car, person, bicycle), and region (e.g., road, sidewalk, vegetation). Anomaly scores are computed using methods tailored to each group’s unique characteristics.
  • For instance, we employ cosine similarity to measure deviations in feature space, as these classes often exhibit distinct geometric and semantic features. The cosine similarity for per-voxel’s logit is computed as:
@@
  • For region, we use entropy-based scoring to capture prediction uncertainty, as these classes typically cover larger, more homogeneous areas. The entropy-based anomaly score is computed as:
@@
    where $\mathcal{K}$ represents the set of possible classes, and $p(k|v_{i})$ denotes the probability that voxel $v_{i}$ is predicted to be of class $k$. To prioritize instance classes, the anomaly scores for region classes are weighted accordingly. Higher entropy values indicate greater uncertainty, which is often associated with OoD regions.
@@
5 Experiments
Two sets of experiments are conducted to evaluate the effectiveness of the proposed OccOoD framework. First, we assessed its performance on OoD on established datasets. Then, we evaluated the performance of semantic occupancy prediction on public benchmarks as well as established datasets. After that, we performed several ablation studies to validate the effectiveness of our design. For additional details on datasets, evaluation metrics, and implementation specifics, please refer to the supplementary materials.
5.1 Results on Different Datasets
Figure 4: Visualization of the results on the VAA-KITTI. From left to right are the input RGB images, the SGN-S [66] results, the OccOoD-S (ours) results, the OoD ground truths, and the semantic occupancy prediction ground truths. Zoom in for a better view.
@@ 5.1 Results on Different Datasets
Quantitative Results: In Table 1, the OoD detection results are reported. Across the three different scales of AuPRCr, our network achieves a significant improvement over the baseline. On the VAA-KITTI dataset, our performance improves by more than $3\%$ in terms of AuPRCr. Similarly, on the VAA-KITTI-360 dataset, our method demonstrates a performance gain of over $5\%$. Furthermore, we present the experimental results of semantic occupancy prediction in Table 2. Our method consistently demonstrates performance improvements over the baseline model, regardless of whether the evaluation is conducted on public datasets or our self-constructed datasets. These results fully validate the effectiveness of our designed VBPF, which progressively integrates features from the BEV perspective, thereby enhancing global context perception. For more details on experimental results, please refer to the supplementary materials.
Qualitative Results: In Fig. 4, four different sets of visualization results are presented. It can be observed that without incorporating a top-down perspective, the baseline network exhibits significant estimation biases. Specifically, by integrating two different observation perspectives, our network achieves more accurate localization for small objects (skunk). Meanwhile, for elongated objects (cow), accurately determining their size from a single image alone is inherently challenging. However, by leveraging multi-view observations, the proposed network makes predictions that are more physically consistent and geometrically plausible.
@@
Table 3: Ablation studies of OccOoD on established datasets.
[…]
Table 4: Ablation and comparison studies of geometry prior on established datasets and regions within 1.2m.
5.2 Ablation Studies
Semantic-aware Anomaly Scoring: Table 3 demonstrates the effectiveness of our proposed Semantic-aware Anomaly Scoring (ASS). Compared to the other three scoring methods, our approach improves AuPRCr by more than $10\%$ across all scales. This is because we incorporate the spatial occupancy characteristics of different objects in 3D space, applying a cosine similarity method for instance-level objects while using an entropy-based approach for larger and more uniform regions. Moreover, in terms of AuROC, we also achieved impressive improvements.
Geometry Prior: In Table 4, we analyze the importance of Geometry Prior by examining AuPRCr and AuROC. With the geometry prior, the network achieves a significant improvement in AuPRCr. In the absence of the geometry prior, the higher AuROC score is due to the fact that it also considers the uncertainty of empty voxel as an anomaly voxel. In our dataset, positive samples are extremely rare, making the AuPRCr metric more suitable for evaluating the model’s performance.
Sequences Used for Training Depth Estimation Model Used for Training Methods #Sequences IoU mIoU Methods Depth Estimation Model IoU mIoU SGN-T [66] 10 46.21 15.32 SGN-S [66] MobileStereoNet [76] 42.76 14.06 OccOoD-T 46.83 16.80 OccOoD-S 43.33 14.54 SGN-T [66] 7 45.15 15.23 SGN-S [66] SPI [45] 41.65 13.68 OccOoD-T 45.49 15.86 OccOoD-S 42.75 13.79
Table 5: Ablation and comparison studies on SemanticKITTI val set [3]. The left table examines different training sequences, whereas the right table studies different depth estimation models.
Impact of Depth Prior: Table 5 demonstrates the impact of different depth models on detection performance. Using only monocular depth estimation, our method outperforms the baseline by $1.1\%$. With the more accurate MSNet3D model [76], the improvement reaches $0.57\%$. The top-down observation perspective mitigates depth map limitations, enabling our network with monocular depth to nearly match the baseline’s performance using stereo-depth maps.
6 Conclusion

### v2

@@ IV-B Overall Architecture
The OccOoD framework, illustrated in Fig. 4, integrates front-view and BEV representations for precise 3D semantic occupancy prediction. 3D voxels preserve fine geometric details, while BEV enhances spatial reasoning and global understanding. To leverage their complementary strengths, we propose the Cross-Space Semantic Refinement (CSSR), which extracts and refines features across both spaces. By fusing multi-scale representations across spaces, CSSR enhances feature representation, sharpens boundaries, and improves OoD detection. Following SGN [36], 2D features extracted from RGB images are transformed into 3D voxel representations via a view transformation module based on MonoScene [9]. These coarse features are refined through a geometry-guided auxiliary 3D occupancy head, supervised by a binary cross-entropy loss $L_{\text{geo}}$, and subsequently filtered by a proposal network trained with an additional binary cross-entropy loss $L_{\text{occ}}$. The filtered features then undergo Geometry Parsing and Semantic Refinement to extract multi-scale BEV features encoding spatial structure and fine-grained semantic voxel features. These BEV features are fused with the original 3D features through a UNet-like module, ensuring precise alignment and enhanced learning. Finally, the upsampled BEV representation is combined with the aggregated voxel features through a 1D convolutional layer to produce the final semantic occupancy prediction.
IV-C Cross-Space Semantic Refinement
The coarse voxel features $F^{3D}$ from the view transformation module and the seed features $F_{\text{s,0}}^{3D}$ from the proposal network are in 3D voxel space. However, relying solely on voxel features limits semantic detail and boundary clarity, harming OoD detection. To address this, we propose Cross-Space Semantic Refinement (CSSR), which leverages voxel and BEV representations to enhance semantic prediction. Features are refined separately in the BEV space to capture global contextual structure and in the voxel space to preserve fine-grained local details and fused to produce accurate occupancy predictions with improved OoD sensitivity.
Geometry Parsing. As shown in Fig. 4, the coarse voxel features $F^{3D}$ are processed through a geometric parsing branch to extract compact BEV representations encoding spatial structure. This branch projects $F^{3D}$ into a lower-dimensional space, applies residual refinement and z-axis alignment, then compresses it via 2D convolutions to generate dense BEV features. These are enhanced by the Vision-RWKV (VRWKV) module [14] through spatial and channel mixing, yielding dense geometric BEV features $\{F_{\text{g}_{i}}^{\text{BEV}}\}_{i}$.
@@ IV-C Cross-Space Semantic Refinement
Semantic Refinement. The seed voxels $F_{\text{s,0}}^{3D}$ from the proposal network undergo semantic refinement to enhance accuracy. The module comprises three cascaded sparse encoder blocks (SEBs), each linked with a BEV projection and a VRWKV block. Each SEB contains a Sparse Feature Encoder (SFE) [49] and a Sparse Geometry Feature Encoder (SGFE) [49], which refine the initial representation through hierarchical pooling and attention, capturing fine-grained details and contextual dependencies.
First, the multi-scale features $F_{\text{s,1}}^{3D}$ and $F_{\text{s,2}}^{3D}$ output by the first two SEBs are projected into BEV space via max-pooling, generating sparse features. These are densified using Spconv and enhanced by the VRWKV module [14] to model long-range dependencies, yielding multi-scale semantic BEV features $\{F_{\text{s}_{i}}^{\text{BEV}}\}_{i}$. The refined voxel features $F_{\text{s}}^{3D}$ are obtained by fusing $F_{\text{s,0}}^{3D}$, $F_{\text{s,1}}^{3D}$, and $F_{\text{s,2}}^{3D}$ via an MLP (Eq. 7):
@@
The semantic refinement stage is optimized using a combined loss of cross-entropy and Lovasz loss [5] as follows:
@@
where $Y_{s}$ represents the semantic labels of seed voxels.
Cross-View Feature Synergy. To integrate BEV and voxel-level representations, we process them in parallel.
For the BEV branch, a UNet-like architecture encodes hierarchical spatial features and decodes high-resolution maps for fusion, including the projected 3D features $F^{\mathrm{BEV}}_{c}$, geometric BEV features $\{F_{\text{g}_{i}}^{\text{BEV}}\}_{i}$, and semantic BEV features $\{F_{\text{s}_{i}}^{\text{BEV}}\}_{i}$. The encoder downsamples $F^{\mathrm{BEV}}_{c}$ through convolutional blocks, each followed by a VRWKV block, and fuses the three BEV streams using an ARF module [37]:
@@
where $F^{\text{BEV}}_{k_{i}}$ and $A^{\text{BEV}}_{k_{i}}$ denote the $k$-th feature map and its attention weight, respectively. The decoder restores resolution via transposed convolutions and skip connections, and the output is reshaped and upsampled to yield a voxel-level representation $\hat{Y}_{b}$.
For the voxel branch, we follow SGN [36], fusing refined semantic features $F_{\text{s}}^{3D}$, non-seed features $F_{\text{n}}^{3D}$, and occupancy features $F^{3D}_{o}$ via Multi-Scale Semantic Propagation (MSSP) as follows:
$\hat{Y}_{\text{$v$}}=\text{MSSP}([F_{\text{s}}^{3D},F_{\text{n}}^{3D},F_{\text{$o$}}^{3D}]).$ (10)
Subsequently, $\hat{Y}_{v}$ and $\hat{Y}_{b}$ are fused via a 1D convolutional layer to produce the final prediction $\hat{Y}_{f}$. The model is optimized with a combined loss as follows:
@@
Loss Function. The total training loss is defined as follows:
@@ IV-D OoD Occupancy Prediction
Geometry Prior. To more efficiently focus on non-empty regions, we introduce a geometry prior by applying an argmax operation to $\hat{Y}_{f}$. This strategy prioritizes potentially anomalous non-empty regions, significantly improving the effectiveness and accuracy of OoD detection.
TABLE I: Quantitative results on the SemanticKITTI validation set [2] and VAA-KITTI. The best and the second best results are in bold and under lined, respectively.
@@
MonoScene [9] 37.12 57.47 27.05 15.72 0.87 14.24 23.55 7.83 0.20 0.77 3.59 18.12 2.57 30.76 1.79 1.03 0.00 6.39 4.11 2.48 11.50
SGN [36] 46.21 59.10 29.41 19.05 0.33 25.17 33.31 6.03 0.61 0.46 9.84 28.93 9.58 38.12 0.47 0.10 0.00 9.96 13.25 7.21 15.32
CGFormer [50] 45.99 65.51 32.31 20.82 0.16 23.52 34.32 19.44 4.61 2.71 7.67 26.93 8.83 39.54 2.38 4.08 0.00 9.20 10.67 7.84 16.89
OccOoD (Ours) 46.83 60.87 32.16 23.52 0.06 26.54 33.51 25.17 0.65 1.00 11.84 28.43 8.90 36.96 0.44 0.16 0.00 9.94 13.00 6.03 16.80
@@
SGN [36] 39.07 46.62 26.54 15.88 0.12 18.34 20.51 0.62 0.44 0.35 1.23 24.76 6.40 21.78 0.48 1.29 0.00 12.25 8.70 4.57 11.10
OccOoD (Ours) 40.04 50.17 27.42 18.40 0.01 17.38 20.19 0.74 0.26 0.56 1.04 25.13 5.53 23.56 0.67 1.95 0.00 11.35 8.15 5.80 11.49
TABLE II: OoD detection performance on VAA-KITTI dataset. Best results in bold.
[…]
Anomalis-Sensitive Scoring. We compute anomaly scores based on the predicted semantic category, employing distinct strategies to capture the two primary forms of uncertainty pertinent to OoD detection.
  • For instance (e.g., car, person), which form compact clusters in the logit space, anomalies appear as outliers from the class centroid. We measure this deviation using cosine similarity:
@@
  • For region (e.g., road, vegetation), which occupy large, homogeneous areas, OoD regions typically induce high prediction uncertainty. This is quantified using entropy:
@@
    where $\mathcal{K}$ represents the set of possible classes, and $p(k|v_{i})$ is the predicted probability of class $k$ for voxel $v_{i}$. To prioritize detection around structured instance classes, the entropy-based scores for region classes are down-weighted. Higher entropy indicates greater uncertainty, a hallmark of OoD regions.
@@
V Experiments
V-A Implementation Details
We train OccOoD (38M parameters) for $35$ epochs on $4$ RTX3090 GPUs with a batch size of $4$, using AdamW with a learning rate of 2e42e{-}4 and weight decay of 1e21e{-}2. We evaluate OccOoD’s semantic occupancy prediction on public benchmarks. For OoD detection, our evaluation is conducted on the proposed datasets with single‑frame settings.
V-B Quantitative Results
Out-of-Distribution Detection. In Table II and III, the OoD detection results are reported. Across the three scales of AuPRCr, our network achieves a significant improvement over the baseline. On the VAA-KITTI dataset, our performance improves by more than $4\%$ in terms of AuPRCr. Similarly, on both the VAA-KITTI-360 and VAA-STU datasets, our method demonstrates superior performance. Notably, CGFormer [50] underperforms in OoD detection despite strong in-distribution accuracy. When encountering anomalies (e.g., animals, garbage bags), it often assigns them to rare in-distribution categories such as bicycle and motorcycle with high confidence, lacking clear semantic correspondence. This overconfident misclassification reflects poor model calibration and an inability to recognize distributional novelty. By enforcing consistency between voxel and BEV spaces, CSSR reduces overconfident misclassification and enhances sensitivity to anomalies.
TABLE III: OoD detection performance on VAA-KITTI-360 and VAA-STU datasets. Best results in bold.
VAA-KITTI-360 VAA-STU
Methods $uparrow AuPRCr\uparrow AuROC\uparrow$
0.8m 1.0m 1.2m 0.8m 1.0m 1.2m
SGN [36] 12.24 24.89 45.15 45.06 2.91 5.74 11.32 64.51
OccOoD (ours) 12.84 25.74 46.28 47.90 3.05 6.15 11.39 65.37
TABLE IV: Quantitative results on the SSCBench-KITTI360 test set [27] and VAA-KITTI-360. The results for counterparts are provided in [27]. The best and the second best results are in bold and under lined, respectively.
Method IoU
car
bicycle
motorcycle
truck
other-veh.
person
road
parking
sidewalk
other-grnd.
building
fence
vegetation
terrain
pole
traf.sign
other-struct.
other-obj.
mIoU
SSCBench-KITTI-360
MonoScene [9] 37.87 19.34 0.43 0.58 8.02 2.03 0.86 48.35 11.38 28.13 3.32 32.89 3.53 26.15 16.75 6.92 5.67 4.20 3.09 12.31
VoxFormer [26] 38.76 17.84 1.16 0.89 4.56 2.06 1.63 47.01 9.67 27.21 2.89 31.18 4.97 28.99 14.69 6.51 6.92 3.79 2.43 11.91
SGN [36] 47.06 29.03 3.43 2.90 10.89 5.20 2.99 58.14 15.04 36.40 4.43 42.02 7.72 38.17 23.22 16.73 16.38 9.93 5.86 18.25
CGFormer [50] 48.07 29.85 3.42 3.96 17.59 6.79 6.63 63.85 17.15 40.72 5.53 42.73 8.22 38.80 24.94 16.24 17.45 10.18 6.77 20.05
OccOoD (Ours) 46.93 28.51 2.95 0.94 9.48 4.57 2.99 58.91 16.17 37.04 5.34 42.99 7.83 37.41 23.01 17.47 18.87 8.94 7.16 18.37
VAA-KITTI-360
SGN [36] 39.90 21.93 0.04 0.00 0.00 0.66 0.88 46.95 13.81 25.67 2.72 33.31 3.07 28.25 12.63 8.95 13.84 3.79 1.52 12.11
OccOoD (Ours) 39.26 22.84 0.00 0.00 0.00 3.54 0.58 49.09 14.65 27.84 3.41 32.59 1.55 27.46 11.46 8.42 12.30 3.63 6.49 12.55
@@ V-B Quantitative Results
3D Occupancy Prediction. We present the experimental results of semantic occupancy prediction in Table I and Table IV. Our method consistently demonstrates performance improvements over the baseline model, regardless of whether the evaluation is conducted on public datasets or our self-constructed datasets. These results validate the effectiveness of our designed CSSR, which leverages complementary strengths of voxel and BEV representations to enhance semantic refinement and global context modeling.
Figure 5: Visualization of the results on VAA-KITTI and VAA-STU datasets. From left to right are the input RGB images, the SGN [36] results, the OccOoD (ours) results, and the OoD ground truths. Zoom in for a better view.
V-C Ablation Studies
TABLE V: Ablation studies of ASS on VAA-KITTI.
Anomaly Score Methods VAA-KITTI
@@ V-C Ablation Studies
0.8m 1.0m 1.2m
Entropy 7.11 12.26 19.95 65.19
Postpro 5.83 9.92 15.94 64.88
Energy 7.24 12.40 20.05 64.78
ASS (Ours) 17.15 23.48 31.83 65.50
Anomalis-Sensitive Scoring. Table V demonstrates the effectiveness of our proposed Anomalis-Sensitive Scoring (ASS). Compared to the other three scoring methods, our approach improves AuPRCr by more than $10\%$ across all scales. This improvement stems from our dual-strategy design, which applies cosine similarity to compact instances and entropy to homogeneous regions, effectively matching the scoring to each object’s spatial characteristics for better OoD identification.
Geometry Prior. Table VI evaluates the role of the geometry prior via AuPRCr and AuROC. Including the prior substantially improves AuPRCr. Without it, the higher AuROC arises because empty voxel uncertainty is also scored as anomalous. Given the scarcity of positive samples in our dataset, AuPRCr is the more appropriate performance metric.
V-D Qualitative Results
Qualitative Comparison. As shown in Fig. 5, unlike SGN [36] which solely relies on voxel input and produces inconsistent shape estimations, our method leverages the complementary strengths of voxel and BEV features through CSSR, effectively improving the localization of small objects (e.g., chair) and maintaining the structural continuity of larger ones (e.g., bear), leading to more complete and geometrically coherent occupancy estimates, ultimately leading to improved OoD detection performance.
TABLE VI: Ablation studies of geometry prior within $1.2m$.
VAA-KITTI VAA-KITTI-360 Methods $uparrow AuPRCr\uparrow AuROC\uparrow$ $w/o$ Geometry prior 14.86 76.59 35.17 36.72 Geometry prior 31.83 65.50 46.28 47.90
Insights Into Real-World Applications. Fig. 6 presents results from a real-world urban environment beyond the VAA-STU dataset. We collected additional real urban scenes using a ZED Mini camera and ran inference on the captured left-view images. The model demonstrates practical efficiency, achieving $9$ FPS on a consumer-grade GPU RTX3090. The left image contains an OoD object (a box on the road). The 3D occupancy prediction (middle) misclassifies it as a motorcycle, exposing a key limitation. In contrast, our proposed OoD detection (right) correctly identifies this anomaly with high confidence, demonstrating the method’s ability to generalize to unseen scenarios and its potential to enhance the safety and reliability of autonomous driving systems.
Figure 6: Results in real-world urban environments.
VI Conclusion